OpenStudy (anonymous):

(1-sec(-x))/(1-cos(-x)) I have to show work, I got this far: (1-secx)/(1+cosx)

OpenStudy (amistre64):

cos(-x) = cos(x)

OpenStudy (anonymous):

Oh!

OpenStudy (amistre64):

other than that, \[\frac{1-\frac{1}{cos}}{1+cos}\] \[\frac{cos-1}{cos(1+cos)}\] \[\frac{cos-(cos^2+sin^2)}{cos+(1-sin^2)}\] \[\frac{cos-cos^2-sin^2}{cos-sin^2+1}\] not really sure what it is you are trying to accomplish so this might be going off into left field

OpenStudy (amistre64):

lol, i got confused and kept the bottom as 1+cos

OpenStudy (anonymous):

None of that showed up, but I have to find the most compact answer.

OpenStudy (amistre64):

sec = 1/cos, sooo 1 - 1/cos --------- 1 - cos and shake it out

OpenStudy (amistre64):

cos - 1 --------- cos(1-cos) cos - 1 --------- -cos(cos-1) -1/cos = -sec

OpenStudy (anonymous):

How did you get that?

OpenStudy (amistre64):

i posted all the work, not sure whats not being followed

OpenStudy (anonymous):

How does 1-1/cos turn into cos-1?

OpenStudy (amistre64):

adding fractions 1 - a/b = (b-a)/b

OpenStudy (anonymous):

Oh, I get that.

OpenStudy (amistre64):

ergo cos-1 1 ----- * ----- cos 1-cos

OpenStudy (anonymous):

What does that go to?

OpenStudy (amistre64):

factor our a -1, and cancel like factors

OpenStudy (anonymous):

And why is (1-cosx) multiplied by cosx?

OpenStudy (amistre64):

(cos-1)/cos = -(1-cos)/cos

OpenStudy (amistre64):

all this is is fractions

OpenStudy (amistre64):

1 - a/b ------- 1-b (1 - a/b) * 1 ------- 1-b b - a 1 ----- * ------- b 1-b

OpenStudy (amistre64):

let b = cos, and a=1

OpenStudy (anonymous):

I get it now, thank you.

OpenStudy (amistre64):

youre welcome